PHYS102 - Gauss’s Law.

Dr. Suess February 2, 2007

PRS Questions slide 2

0.1 Flux - General

Question #1

Consider the Gaussian surface of a cube which encloses a charge of +q that is located in the exact middle of the cube. What is the electric flux through one side of the cube?

1. 0.

2. q/ε_{0}

3. q/6 ε_{0}

4. There is enough information, but none of the above are correct.

5. There is not enough information.

PHYS102 Gauss’s Law - Conductors – slide 2

Answer to Question #1

• The electric flux is given by Φ = \frac{Q_{enclosed}}{ε_{0}} .

• Q is the amount of charge contained inside the closed surface (in this case Q_{enclosed} = +q).

• Since the charge q is located in the center of the cube, then each face of the cube will have the same number of field lines passing through its surface.

• Each face will contribute the same of amount of electric flux (and there are 6 faces to the cube).

• The answer is 3.

PHYS102 Gauss’s Law - Conductors – slide 3